# Mean-field limits for Consensus-Based Optimization and Sampling

### v2

@@ Conclusion
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{2(q-p)}{q}}\right),$
which concludes the proof. ∎
@@ 5.2 Handling the case where $u>\ell>0$ in Assumption H
Our main results stand only when $f\in\mathcal{A}(s,\ell,u)$ with $u=\ell$. This is a convenient setting because it then holds by Proposition A.3 that the $p$-th moment of the reweighted measure $L_{\beta}\mu$ is controlled by the $p$-th moment of the $\mu$, which is key in order to prove well-posedness of the mean-field system, to obtain moment estimates for empirical measures, and, most importantly, to improve the Wasserstein stability estimates for the weighted mean and weighted covariance, see Corollary 3.3. In the interest of simplicity, we refrained from stating precise results for the cases $u>\ell=0$ and $u>\ell>0$. The former remains open, but we note that a quantitative mean-field estimate can still be obtained in the latter setting. The main ingredients to this end are the following:
  • Moment estimates for the empirical measure and well-posedness of the mean-field equation can be obtained by using Proposition A.3 together with the Bihari–LaSalle inequality instead of Grönwall’s inequality, using the same reasoning as in Lemma 3.5 and Section 4, respectively. In contrast with the setting $\ell=u$, these results hold only up to some finite time $T<T_{+}$, where $T_{+}\in(0,\infty)$ is the time at which the bounds through the Bihari–LaSalle inequality become singular.
  • In order to obtain a mean-field result up to some time $T<T_{+}$, the approach presented in Subsection 5.1 can be used, this time with the stopping time $\theta_{J}:=\tau_{J}\wedge\overline{\tau}_{J}$, where
@@
    Employing this stopping time instead of (5.2) allows for using the basic stability estimates for the weighted moments of Proposition 3.1, which are valid even when $\ell<u$, instead of the improved stability estimates in Corollary 3.3, which require that $\ell=u$. In order to complete the proof as in Subsection 5.1, it remains to bound the probability $\mathbf{P}\left[\theta_{J}\leqslant T\right]$, which can be achieved by noticing that
@@
    The first term is bounded from above by $CJ^{-2}$ by Lemma 2.5, whereas the second term can be bounded by $CJ^{-\frac{p}{2}}$ as in the proof in Subsection 5.1. Thus, the probability $\mathbf{P}[\theta_{J}(R)\leqslant T]$ scales as $J^{-\min\{2,\frac{p}{2}\}}$, instead of $J^{-2}$ in the case where $u=\ell$. As a consequence, we obtain a mean-field result of the type
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\min\left\{2,\frac{p}{2}\right\}\frac{q-p}{q}}.$
@@ A.1 Stability estimates for ratios of integrals
$\quad+\left\lVert{g(y)\mathcal{S}(\nu)}\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\right\rVert_{L^{2}(\pi)}.$
@@
$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}W_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R.$ (A.4)
@@ A.2 Bound on weighted moments
$\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x).$

### v3

@@ Conclusion
$\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant C\left(J^{-\frac{p}{2}}+J^{-\frac{\alpha(q-p)}{q}}\right),$
which concludes the proof since $q$ and $\alpha$ can be taken arbitrarily large. ∎
@@ 5.2 Handling the case where $u>\ell>0$ in Assumption H
Our main results stand only when $f\in\mathcal{A}(s,\ell,u)$ with $u=\ell$. This is a convenient setting because it then holds by Proposition A.3 that the $p$-th moment of the reweighted measure $L_{\beta}\mu$ is controlled by the $p$-th moment of $\mu$, which is key in order to prove well-posedness of the mean-field system, to obtain moment estimates for empirical measures, and, most importantly, to improve the Wasserstein stability estimates for the weighted mean and weighted covariance, see Corollary 3.3. In the interest of simplicity, we refrained from stating precise results for the cases $u>\ell=0$ and $u>\ell>0$. We do not address the former setting, but note that a quantitative mean-field estimate can still be obtained in the latter. The main ingredients to this end are the following:
  • Moment estimates for the empirical measure and well-posedness of the mean-field equation can be obtained by using Proposition A.3 together with the Bihari–LaSalle inequality instead of Grönwall’s inequality, using the same reasoning as in Lemma 3.5 and Section 4, respectively. In contrast with the setting $\ell=u$, these results hold only up to some finite time $T<T_{+}$, where $T_{+}\in(0,\infty)$ is the time at which the bounds given by the Bihari–LaSalle inequality become singular.
  • In order to obtain a mean-field result up to some time $T<T_{+}$, the approach presented in Subsection 5.1 can be used, but with the stopping time $\theta_{J}:=\tau_{J}\wedge\overline{\tau}_{J}$, where
@@
    Employing this stopping time instead of (5.2) enables to use the basic stability estimates for the weighted moments of Proposition 3.1, which are valid even when $\ell<u$, instead of the improved stability estimates in Corollary 3.3, which require that $\ell=u$. It is then simple to obtain a propagation of chaos estimate for the stopped particle system, of the form
    In order to complete the proof as in Subsection 5.1, it remains to bound the probability $\mathbf{P}\left[\theta_{J}\leqslant T\right]$, which can be achieved by noticing that
@@
    By Lemma 2.5, For any $\alpha>0$ there exists $C_{\alpha}>0$ such that the first term is bounded from above by $C_{\alpha}J^{-\alpha}$ for all $J\in\mathbf{N}^{+}$, whereas the second term can be bounded by $CJ^{-\frac{p}{2}}$ as in the proof presented in Subsection 5.1. Thus, the probability $\mathbf{P}[\theta_{J}(R)\leqslant T]$ scales as $J^{-\min\left\{\alpha,\frac{p}{2}\right\}}$, instead of $J^{-\alpha}$ in the case where $u=\ell$. As a consequence, we obtain a mean-field result of the type
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\frac{p}{2}\left(\frac{q-p}{q}\right)}.$
    Since $q$ was arbitrary given the assumption that $\overline{\rho}_{0}$ has infinitely many moments, it follows that for every $\varepsilon>0$, there exists $C_{\varepsilon}$ such that
    $\forall J\in\mathbf{N}^{+},\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X_{t}^{(j)}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\frac{p}{2}\left(1-\varepsilon\right)}.$
    Hence, we recover almost the classical Monte Carlo rate. By using the same method as in [vaes2024sharppropagationchaosensemble], it is in fact possible to prove that this statement holds also for $\varepsilon=0$, but for conciseness we do not present the details here.
@@ A.1 Stability estimates for ratios of integrals
$\quad+\left\lVert{g(y)\mathcal{S}(\nu)}\left(\frac{1}{\sqrt{\mu[h]}}-\frac{1}{\sqrt{\nu[h]}}\right)\right\rVert_{L^{2}(\pi)},$
@@
$\forall\mu\in\mathcal{P}_{p,R}(^{d}),\qquad\big{\lvert}\mathcal{S}(\mu)\big{\rvert}\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}W_{p}(\delta_{0},\mu)\leqslant\big{\lvert}\mathcal{S}(\delta_{0})\big{\rvert}+L_{\mathcal{S}}R,$ (A.4)
@@ A.2 Bound on weighted moments
$\forall\mu\in\mathcal{P}_{p}(\mathbf{R}^{d}),\qquad\frac{\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}{\int_{\mathbf{R}^{d}}\operatorname{e}^{-\beta f(x)}\,\mu(\mathrm{d}x)}\leqslant C\int_{\mathbf{R}^{d}}\lvert x\rvert^{p}\,\mu(\mathrm{d}x)$